Roger is 6 years older than Zoe. Three years ago, Zoe's age was two-thirds of Roger's age. Find their present ages.
*Let Roger's age be x and Zoe's age be y

Answer by mananth(16946)

Zoe's age be x
Roger will be x+6
..
3 years ago
Zoe was x-3
Roger was x+6-3=x+3 years
x-3 = 2/3 * x+3
3(x-3)=2(x+3)
3x-9=2x+6
x=15 that is Zoes's age
Roger is 15+6 = 21 years